# Application of Non-Linear Noise Regression in the Virgo Detector

### v3

@@ 3.2 Training and cleaning
There are a number of hyperparameters which can optimize the performance of DeepClean, including the learning rate, kernel size, stride etc. For all of these, we use the same settings as used in Saleem et al. 2023. The data from all the witness channels are first pre-processed such that each timeseries has zero mean and unit variance. These normalized datasets are bandpassed and then divided into overlapping kernels  [^4] The kernels are here the segments of data and are distinct from the convolution filter kernels in CNN architecture that carry the weights.. with kernel size of 8s, and 7.75 s overlaps. The 4096 s of training data yields 16353 overlapping kernels, calculated using the formula:
@@
The cleaned strain is produced in chunks of 4096  s for all three bands. In addition to the normalization of the input witness channels mentioned earlier, there are batch normalizations applied at each layer of the neural network to ensure consistent input distributions for the subsequent layers. In DeepClean, batch normalization also makes sure that the final noise predictions have nearly zero mean and unit variance. For this reason the predicted noise is subtracted from normalized version of the raw strain, which yields cleaned strain in normalized units. This is then un-normalized with the mean and standard deviation used earlier to normalize the unclean strain. Following that, the cleaned strain is bandpassed to the target frequency band, in order to make sure that DeepClean did not make any noise prediction outside the desired frequency band. Hereafter in this paper, cleaned strain refers to this post-processed cleaned strain.
4 Results
4.1 Improvement in the amplitude spectral density (ASD)
4.1.1 Single layer subtractions and their ASD ratio
The primary approach used here for assessing DeepClean’s performance is the comparison of the ASD of the cleaned strain (labeled V1:DC) against the original unclean strain (labeled V1:ORG), where V1 denotes the Virgo detector. Figure 3 illustrates the ASD enhancements for the targeted frequency ranges of 98–110 Hz, 142–162 Hz, and 197–208 Hz. The upper panels in each set of plots display the ASDs, while the lower panels show the ASD ratios of the cleaned strain to the uncleaned strain. Significant subtractions are observed across all three bands. The efficacy of DeepClean varied significantly with the length of the training data. Although segments of 1024 s and 2048 s were used, the results were suboptimal compared to those obtained from the 4096 s segments. Figure 3 is produced using a 4096 s segment.
This may be attributed to the non-stationary features in the data during O3b. The data set is characterized by periods in which the auxiliary sensors exhibit high coherence with the readout channel for durations as brief as 200 s, followed by incoherent intervals. Longer training segments, such as 4096 s, are advantageous, as they encompass a broader spectrum of these variable conditions, thereby improving the model’s ability to accurately predict and mitigate them. Consequently, training with 4096 s segments proved to be more effective for O3b Virgo data.
@@ 4.1.1 Single layer subtractions and their ASD ratio
4.1.2 Multi-training and ASD ratio
@@
4.2 Improvement of BNS inspiral range
The application of the DeepClean algorithm for noise subtraction in GW detectors yields quantifiable improvements in the BNS inspiral range, which is a measure of the average distance at which a binary neutron star (BNS) system with component masses of $m_{1}=m_{2}=1.4\,M_{\odot}$ can be detected with a SNR of 8. This range effectively serves as an indicator of overall GW detector sensitivity (Finn & Chernoff, 1993; Chen et al., 2021; Finn, 1996). An extended BNS inspiral range reflects heightened detector sensitivity, a direct result of advanced noise reduction methods that enable the capture of fainter, more distant signals.
4.2.1 BNS range from single layer
@@
This BNS inspiral range shows a daily fluctuation, with maximal values occurring around midnight and at minimal values after 9:00 AM (UTC), driven by local environmental noise sources and site-specific dynamics. These fluctuations primarily stem from changes in anthropogenic noise (e.g., traffic vibrations) and wind conditions (Coughlin et al., 2011; Fiori et al., 2020). Nighttime brings optimal sensitivity due to minimal human activity and lower seismic noise, while morning traffic increases ground vibrations, reducing sensitivity.
@@
Figure 5: Analysis of the inspiral range sensitivity for BNS, using Virgo O3b data from GPS time 1265127585 to 1265352161. The target frequency band for this analysis is the 142–162 Hz range. This plot highlights the DeepClean effectiveness in enhancing BNS inspiral detection over each 4096 s segment, represented by the “Observation Index.” Note: The periodic fluctuations reflect anthropogenic noise near the Virgo detector, with the highest sensitivity around midnight and the lowest around 09:00 (UTC), when anthropogenic noise is most prevalent.
4.2.2 BNS range from multi-training
@@
4.3 Performance evaluation using CBC injections
@@
We now conduct an injection analysis using CBC signals. The purpose is to verify that the denoising with DeepClean does not alter the GW signals that are present in the data, and also to examine any improvements in the credible intervals of the estimated parameters (Christensen & Meyer, 2022).
4.3.1 GW generation and event selection
@@
In the following, we use utilities LALSuite and Bilby libraries (LIGO Scientific Collaboration et al., 2018; Ashton et al., 2019; Wette, 2020) to generate Binary Black hole (BBH) signals. The component masses $m_{1}$ and $m_{2}$ are drawn from a uniform distribution, ensuring that the total mass $M$ corresponds to innermost stable circular orbit (ISCO) frequencies $f_{\text{ISCO}}$ falling within the three targeted frequency bands described above. A similar criterion was employed also by Saleem et al. 2023 for distributing the injection parameters.
@@
4.3.2 Signal-to-noise ratio analysis
Following the procedure described in Section 4.3.1, we inject a total of 3,200 GW signals from BBH mergers into the dataset, spanning the full frequency range of 15–415 Hz. The SNR of each injected signal was computed using the matched-filter function provided by PyCBC (Biwer et al., 2019), with a waveform template corresponding to the injected signal.
For each injection, the matched-filter SNR was calculated both before and after applying DeepClean.
@@ 4.3.2 Signal-to-noise ratio analysis
To analyze these results further, we categorized the SNR fractional differences into three groups based on a threshold: unchanged ($\lvert\frac{\Delta\mathrm{SNR}}{\mathrm{SNR}}\rvert\leq 1\%$), increased ($\frac{\Delta\mathrm{SNR}}{\mathrm{SNR}}>1\%$), and decreased ($\frac{\Delta\mathrm{SNR}}{\mathrm{SNR}}<-1\%$). For multi-band subtraction, 71.3% of GW signals showed an increase, 28.5% showed a decrease, and only 0.2% remained unchanged. For the three single bands–98–110 Hz, 142–162 Hz, and 197–208 Hz–the percentages of injections with SNR gains were 74.8%, 45.3%, and 48.8%, respectively. The remainder in each band either experienced a decrease or remained unchanged (with only 0.2–3.2% unchanged per band).
@@
4.3.3 Parameter Estimation: Testing DeepClean’s Impact on GW Signal structure
In this section, we describe a parameter estimation study to confirm that DeepClean does not adversely affect the astrophysical signals present in the data. We injected 128 BBH coalescence signals with ISCO frequencies ranging from 15 to 415 Hz, over a 4096 s interval starting at GPS time 1265127585. For a BBH signal, at least 15 independent parameters need to be estimated simultaneously. In this study, we adopt a one-at-a-time approach, estimating a single parameter of interest while fixing all other parameters to their true or injected values. This approach allows us to isolate and analyze the effects on individual parameters more effectively.
A parameter estimation validation study typically serves two purposes. The first and most critical goal is to test whether the signal retains its original structure after noise regression, which would manifest as biases in the estimated parameters. The second goal is to evaluate any improvement in the uncertainty of the estimated parameters resulting from noise removal. For the former, the one-at-a-time approach is particularly suitable because the full 15-dimensional parameter space, with its inherent degeneracies, may mask subtle biases by redistributing them among degenerate parameters. For example, if the bandpass filters used in DeepClean introduce additional phase to the signal, this would bias the coalescence phase parameter. However, such a bias could be reabsorbed into other parameters, such as the time of arrival or the sky location of the source. By fixing all other parameters to their true values, we eliminate the possibility of such reabsorption, ensuring that any additional phase is directly captured as a bias. Without the one-at-a-time approach, a full 15-dimensional parameter estimation would be necessary to obtain realistic uncertainty estimates. Since the focus of this study is on preserving the signal structure, we exclusively rely on the one-at-a-time method, which is also computationally much less demanding.
@@ 4.3.3 Parameter Estimation: Testing DeepClean’s Impact on GW Signal structure
In Figure 9, we present example posterior distributions for the chirp mass $\mathcal{M}c$, mass ratio $q$, component spin magnitudes $a_{1}$ and $a_{2}$, inclination angle $\theta_{jn}$, and luminosity distance $d_{L}$ for one representative injection (injection 24). The blue solid and orange dashed curves represent the posterior distributions obtained from the uncleaned and DeepClean strains, respectively, while vertical dashed lines indicate the injected (true) parameter values. We observe that the posterior peaks derived from the uncleaned strain exhibit slight offsets from the injected values due to the random nature of noise. In contrast, posteriors from the DeepClean strain align more accurately with both the injected and zero-noise values, demonstrating reduced bias. For comparison, the green dotted curves represent posteriors obtained from the same injection but without noise.
@@
However, the PP plots from the uncleaned data show clear deviations, notably falling below the diagonal for parameters such as $\theta_{jn}$, and $d_{L}$. These deviations indicate biases caused by non-Gaussian noise. These findings underscore that the application of DeepClean with a multi-training approach not only efficiently reduces residual noise over a wide frequency range but also preserves gravitational-wave signal integrity and decreases uncertainties in parameter measurements. These frequency-sensitive parameters also show slight reductions in posterior uncertainties, as illustrated by the narrower posterior distributions in Figure 9.
@@
Figure 10: The PP plots of the six parameters with unclean and DeepClean strain showing that DeepClean does not affect the recovery of the underlying physical parameters.
5 Summary and Conclusions
One of the main challenges faced by noise regression pipelines such as DeepClean is the large number of witness channels available, many of which exhibit non-negligible coherence with the strain channel. Handling more than a few tens of channels exceeds our computational capacity, even with state-of-the-art resources (Reissel et al., 2025). In this paper, we demonstrate that identifying and dividing the channels into smaller frequency bands achieves significant success in noise subtraction. DeepClean, previously demonstrated with the two LIGO detectors, showcases its capability for non-linear noise regression in Virgo, a detector with distinct design features, thereby highlighting its generality. Additionally, we employ a computationally efficient alternative for validating the subtraction by using a one-at-a-time parameter estimation approach.
In this study of the DeepClean framework applied to the Virgo detector data, we demonstrated an increase in sensitivity of the Virgo detector, improving the detection prospects for GW signals. The results are promising and we expect the method can be used in the ongoing O4b campaign for noise reduction. In the future, there are a number of challenges we plan to tackle:
  • Online analysis: All analysis in this paper was performed offline, however, online processing will be critical to take advantage of the noise improvemets in O5 , enhancing real-time noise reduction in the Virgo detector and improving pre-merger GW signal detection.
  • High computational demands: Analyzing Virgo data to capture the 150 Hz peak and its sidebands via ASC witness sensors required significant computational resources, often compromising between data depth and computing power, and we will explore alternative architectures that may be more memory efficient.

### v4

@@ III.2 Training and cleaning
There are a number of hyperparameters which can optimize the performance of DeepClean, including the learning rate, kernel size, stride etc. For all of these, we use the same settings as used in [19]. The data from all the witness channels are first pre-processed such that each timeseries has zero mean and unit variance. These normalized datasets are bandpassed and then divided into overlapping kernels  [^4] The kernels are here the segments of data and are distinct from the convolution filter kernels in CNN architecture that carry the weights.. with kernel size of 8s, and 7.75 s overlaps. The 4096 s of training data yields 16353 overlapping kernels, calculated using the formula:
@@
The cleaned strain is produced in chunks of 4096  s for all three bands. In addition to the normalization of the input witness channels mentioned earlier, there are batch normalizations applied at each layer of the neural network to ensure consistent input distributions for the subsequent layers. In DeepClean, batch normalization also makes sure that the final noise predictions have nearly zero mean and unit variance. For this reason the predicted noise is subtracted from normalized version of the raw strain, which yields cleaned strain in normalized units. This is then un-normalized with the mean and standard deviation used earlier to normalize the unclean strain. Following that, the cleaned strain is bandpassed to the target frequency band, in order to make sure that DeepClean did not make any noise prediction outside the desired frequency band. Hereafter in this paper, cleaned strain refers to this post-processed cleaned strain. Table 1 presents the configuration used for training and cleaning.
Table 1: Summary of the DeepClean training configuration with Virgo O3b data. The full configuration files are publicly available at  [^6] https://github.com/weizmannk/Virgo-DeepClean/blob/main/config/config_virgo.ini.
Parameter Value
Data Segmentation
Sampling frequency 4096 Hz
Length of overlapping kernels 8 s
Stride (training / cleaning) 0.25 s / 4 s
Pre-processing
Bandpass filter (8th order, Hanning window,
median padding)
Training Settings
Optimizer Adam
Batch size 32
Epochs 30
Learning rate $10^{-3}$
Weight decay $10^{-5}$
Training duration 4096 s
Training cadence $10^{5}$ s
At-a-time cleaning duration 4096 s
Dataset
Detector run Virgo O3b
Training/validation data split 90% (10% validation)
III.3 Performance evaluation with amplitude spectral density
III.3.1 Single layer subtractions and their ASD ratio
The primary approach used here for assessing DeepClean’s performance is the comparison of the amplitude spectral density (ASD) of the cleaned strain (labeled V1:DC) against the original unclean strain (labeled V1:ORG), where V1 denotes the Virgo detector. Figure 3 illustrates the ASD enhancements for the targeted frequency ranges of 98–110 Hz, 142–162 Hz, and 197–208 Hz. The upper panels in each set of plots display the ASDs, while the lower panels show the ASD ratios of the cleaned strain to the uncleaned strain. Significant subtractions are observed across all three bands. The efficacy of DeepClean varied significantly with the length of the training data. Although segments of 1024 s and 2048 s were used, the results were suboptimal compared to those obtained from the 4096 s segments.
This may be attributed to instabilities present in some witness channels during O3b, where sensors intermittently failed to record data, resulting in periods of low or zero coherence with the strain channel. Shorter training segments are more likely to overfit to these unstable intervals, while longer segments promote learning of persistent coupling patterns. This is consistent with the improved subtraction performance observed for 4096 s training compared to shorter durations. Specifically, Figure 3 shows that the model trained with 4096 s segments achieves more than 99.9% suppression at the dominant narrow peaks (top panel), more than 88% suppression across all peaks (bottom panel), and a median reduction of about 72% in the 149–152 Hz band ( middle panel).
@@ III.3.1 Single layer subtractions and their ASD ratio
III.3.2 Multi-training and ASD ratio
@@
III.4 Performance evaluation with BNS inspiral range
The application of the DeepClean algorithm for noise subtraction in GW detectors yields quantifiable improvements in the BNS inspiral range, which is a measure of the average distance at which a binary neutron star (BNS) system with component masses of $m_{1}=m_{2}=1.4\,M_{\odot}$ can be detected with a SNR of 8. This range effectively serves as an indicator of overall GW detector sensitivity [33, 34, 35]. An extended BNS inspiral range reflects heightened detector sensitivity, a direct result of advanced noise reduction methods that enable the capture of fainter, more distant signals.
III.4.1 BNS range from single layer
@@
This BNS inspiral range shows a daily fluctuation, with maximal values occurring around midnight and at minimal values after 9:00 AM (UTC), driven by local environmental noise sources and site-specific dynamics. These fluctuations primarily stem from changes in anthropogenic noise (e.g., traffic vibrations) and wind conditions [36, 37]. Nighttime brings optimal sensitivity due to minimal human activity and lower seismic noise, while morning traffic increases ground vibrations, reducing sensitivity.
@@
Figure 5: Analysis of the inspiral range sensitivity for BNS, using Virgo O3b data from GPS time 1265127585 to 1265352161. The target frequency band for this analysis is the 142–162 Hz range. This plot highlights the DeepClean effectiveness in enhancing BNS inspiral detection over each 4096 s segment, represented by the “Observation Index.” Note: The periodic fluctuations reflect anthropogenic noise near the Virgo detector, with the highest sensitivity around midnight and the lowest around 09:$0$ (UTC), when anthropogenic noise is most prevalent.
III.4.2 BNS range from multi-training
@@
III.5 On the nature of the subtracted noise: linear vs non-linear
To determine whether the Virgo noise subtracted by DeepClean includes non-linear couplings, we designed a dedicated test targeting the role of non-linear activation functions. Specifically, we trained two versions of the same network architecture using all 122 witness channels in the 150 Hz subtraction band: one with the standard tanh activation functions after each convolution and batch normalization layer (refer to Figure 1), and another with these activations removed. While the resulting network without tanh is not strictly linear, it is significantly restricted in its ability to model non-linear couplings.
We show the results in Figure 7, left panel, where both models were able to reduce the central peak at 150 Hz, which is known to be a linearly coupled feature (with the model without activations performing slightly better). However, only the model with non-linear activation functions effectively removed the adjacent sidebands, suggesting that their subtraction required modeling non-linear couplings between witness channels. This result does not represent a comprehensive classification of the linear or non-linear nature of the Virgo noise, but it provides strong evidence that at least some of the subtracted noise is non-linear. More importantly, it demonstrates that with the existing set of Virgo witness channels, DeepClean is capable of learning and removing non-linear noise couplings.
Beyond that, there are practical advantages to using DeepClean-like tools even in the case of linear noise. In a separate study, we tested two training strategies using the same 122 witness channels: (1) single-pass training, where all channels are used together in a single training session; and (2) multi-layer training, where the channels are divided into eight groups containing 14, 32, 32, 12, 8, 5, 9, and 10 channels, respectively. These groups were formed based on channel types—such as magnetic, electric, seismic, and alignment sensing and control—and processed sequentially. In the multi-group strategy, the cleaned strain from one group is passed forward to the next. The results, shown in Figure 7, right panel, indicate that the single-pass strategy consistently outperforms the multi-layer approach. This demonstrates that multi-layer subtraction remains sub-optimal, even when each group contains many channels.
The sub-optimality arises because the optimization within each group is local and does not account for inter-group dependencies. That is, the coupling coefficients for one group are learned assuming fixed values (zero) for the coefficients associated with the remaining channels, thereby restricting the optimization to a local minimum in the loss function. In contrast, the single-pass strategy performs a joint optimization over the full parameter space, increasing the likelihood of reaching a global minimum.
This illustrates the ability of deep learning tools like DeepClean to perform optimal regression over high-dimensional input spaces, enabled by hardware-accelerated training on GPUs. Traditional analytical linear methods, such as Wiener filtering, become increasingly difficult to scale due to computational and numerical limitations when many channels are involved. Deep learning frameworks allow all relevant channels to be used simultaneously, enhancing performance regardless of whether the underlying couplings are linear or non-linear.
Figure 7: Left: Comparison of the ASD and ASD ratios between the original Virgo data (V1:ORG) and the data processed with two versions of the DeepClean network: one with standard non-linear tanh activations (red), and one with activations turned off (purple). Both models reduce the 150 Hz line, a linearly coupled feature, but only the non-linear model reduces the nearby sidebands—suggesting that their subtraction requires modeling non-linear noise couplings. Right: Comparison between two training strategies for DeepClean using all 122 witness channels: single-pass (red), which jointly optimizes over all channels in a single training session, and multi-layer (purple), where the channels are divided into eight groups and processed sequentially. The single-pass model consistently outperforms the multi-layer strategy, highlighting the importance of joint optimization across all channels to capture inter-channel dependencies and avoid local minima. The bottom panel shows the corresponding ASD ratios relative to the original data.
IV Validation analysis using CBC injections
@@
We now conduct an injection analysis using CBC signals. The purpose is to verify that the denoising with DeepClean does not alter the GW signals that are present in the data, and also to examine any improvements in the credible intervals of the estimated parameters [38].
IV.1 GW generation and event selection
@@
In the following, we use utilities LALSuite and Bilby libraries [39, 40, 41] to generate binary black hole (BBH) signals. The component masses $m_{1}$ and $m_{2}$ are drawn from a uniform distribution, ensuring that the total mass $M$ corresponds to innermost stable circular orbit (ISCO) frequencies $f_{\text{ISCO}}$ falling within the three targeted frequency bands described above. A similar criterion was employed also by [19] for distributing the injection parameters.
@@
IV.2 Signal-to-noise ratio analysis
Following the procedure described in Section IV.1, we inject a total of 3,200 GW signals from BBH mergers into the dataset, spanning the full frequency range of 15–415 Hz. The SNR of each injected signal was computed using the matched-filter function provided by PyCBC [42], with a waveform template corresponding to the injected signal. For each injection, the matched-filter SNR was calculated both before and after applying DeepClean. In both cases, the filtering step used the same injected waveforms as templates, ensuring a direct comparison.
@@ IV.2 Signal-to-noise ratio analysis
To analyze these results further, we categorized the SNR fractional differences into three groups based on a threshold: unchanged ($\lvert\frac{\Delta\mathrm{SNR}}{\mathrm{SNR}}\rvert\leq 1\%$), increased ($\frac{\Delta\mathrm{SNR}}{\mathrm{SNR}}>1\%$), and decreased ($\frac{\Delta\mathrm{SNR}}{\mathrm{SNR}}<-1\%$). For multi-band subtraction, 71.3% of GW signals showed an increase, 28.5% showed a decrease, and only 0.2% remained unchanged. For the three single bands–98–110 Hz, 142–162 Hz, and 197–208 Hz–the percentages of injections with SNR gains were 74.8%, 45.3%, and 48.8%, respectively. The remainder in each band either experienced a decrease or remained unchanged (with only 0.2–3.2% unchanged per band). This test demonstrates that signal recovery by search pipeline remains robust after the application of DeepClean. Even though the gain in SNR is not substantial, we observe a slightly higher number of sources with improved SNR compared to those with reduced SNR.
@@
IV.3 Parameter Estimation: Testing DeepClean’s Impact on GW Signal structure
To evaluate whether DeepClean affects the recovery of astrophysical signals, we perform a set of parameter estimation (PE) studies using injected binary black hole (BBH) signals. We adopt two complementary strategies: a one-at-a-time approach, where individual parameters are varied while others are fixed at their true injected values, and a full-dimensional analysis, where all 15 BBH parameters are simultaneously estimated.
Each method has distinct advantages. The one-at-a-time approach allows us to isolate the behavior of specific parameters and is particularly useful for detecting subtle biases that might otherwise be absorbed by degeneracies in the full parameter space. For example, if DeepClean introduces additional phase shifts, the effect on the coalescence phase may be redistributed among correlated parameters like time or sky location in a full-dimensional fit. Fixing other parameters ensures that such distortions are directly attributed to the parameter of interest. This method is also computationally efficient and ideal for detecting structural distortions.
In contrast, simultaneous estimation across all 15 parameters yields more realistic posterior uncertainties and provides a holistic view of how DeepClean affects parameter recovery in practice. However, it may obscure small systematic effects due to parameter correlations. Additionally, fixing all but one parameter in the one-at-a-time method may overly constrain the waveform space, preventing the model from exploring alternative morphologies that could better fit a subtly distorted signal. Since neither method is universally more sensitive—each depending on the nature of potential distortions—we perform both to obtain a comprehensive validation.
In the following, we present the analysis and results from both approaches.
IV.3.1 One-at-a-Time Parameter Estimation
To assess whether DeepClean preserves the structure of astrophysical signals, we perform a validation study using the one-at-a-time parameter estimation approach. We injected 128 BBH coalescence signals with ISCO frequencies ranging from 15 to 415,Hz into a 4096,s segment of data, starting at GPS time 1265127585. For each injection, we estimate a single parameter of interest while fixing all others to their true (injected) values, allowing us to isolate DeepClean’s impact on that parameter.
@@ IV.3.1 One-at-a-Time Parameter Estimation
In Figure 10, we present example posterior distributions for the chirp mass $\mathcal{M}c$, mass ratio $q$, component spin magnitudes $a_{1}$ and $a_{2}$, inclination angle $\theta_{jn}$, and luminosity distance $d_{L}$ for one representative injection (injection 24). The blue solid and orange dashed curves represent the posterior distributions obtained from the uncleaned and DeepClean strains, respectively, while vertical dashed lines indicate the injected (true) parameter values. We observe that the posterior peaks derived from the uncleaned strain exhibit slight offsets from the injected values due to the random nature of noise. In contrast, posteriors from the DeepClean strain align more accurately with both the injected and zero-noise values, demonstrating reduced bias. For comparison, the green dotted curves represent posteriors obtained from the same injection but without any noise realization.
@@
However, the PP plots from the uncleaned data show clear deviations, notably falling below the diagonal for parameters such as $\theta_{jn}$, and $d_{L}$. These deviations indicate biases caused by non-Gaussian noise. These findings underscore that the application of DeepClean with a multi-training approach not only efficiently reduces residual noise over a wide frequency range but also preserves GW signal integrity and decreases uncertainties in parameter measurements. These frequency-sensitive parameters also show slight reductions in posterior uncertainties, as illustrated by the narrower posterior distributions in Figure 10.
@@
Figure 11: The PP plots of the six parameters with unclean and DeepClean strain showing that DeepClean does not affect the recovery of the underlying physical parameters. The shaded regions represent expected statistical fluctuations (68%, 95%, and 99.7% confidence intervals). Combined Kolmogorov–Smirnov p-values are $\simeq 3.47\times 10^{-14}$ (Unclean) and $\simeq 0.0212$ (DeepClean). The largest pre-cleaning deviations are for $\theta_{jn}$ ($5.45\times 10^{-7}$) and $d_{L}$ ($1.46\times 10^{-6}$); a milder deviation is seen for the mass ratio $q$ ($0.0275$). After DeepClean, these improve to $\simeq 0.0371$, $0.00585$, and $0.679$, respectively, while the chirp mass $M_{c}$ is near-uniform in both cases ($0.361\to 0.54$).
IV.3.2 Simultaneous Estimation of all 15 Parameters
In this section, we present a parameter estimation (PE) study where all 15 independent parameters are estimated simultaneously.
We consider the same injections as in the one-at-a-time analysis but analyse only 60 signals due to computational reasons. Note that this is a single-detector analysis, and hence it is difficult to constrain extrinsic parameters such as time and phase of arrival, luminosity distance, and the angles describing the source’s location and orientation. In Figure 12, we show the PP plots for chirp mass, mass ratio, the two spin magnitudes, inclination angle, and luminosity distance. As expected, the PP plots for inclination angle and luminosity distance show poor recovery of the true values, mostly reflecting the priors. However, for masses and spins, the PP plots indicate good and comparable recovery both before and after applying DeepClean. There is no evidence of any bias introduced by the application of DeepClean.
Figure 12: PP plots for six parameters, comparing uncleaned data and data processed with DeepClean, using full 15-dimensional parameter estimation. The shaded regions represent expected statistical fluctuations (68%, 95%, and 99.7% confidence intervals). The results indicate that DeepClean preserves the integrity of the GW signal for all parameters except the luminosity distance ($d_{L}$), which shows a systematic deviation likely due to inherent parameter degeneracies. The combined Kolmogorov–Smirnov p-values are $\simeq 2\times 10^{-4}$ for the Unclean data and $\simeq 1\times 10^{-4}$ after DeepClean
While the PP plot can be used as a measure of credibility of the injection recovery, it alone does not show whether the noise subtraction has improved the error bars. For that matter, in Figure 13, we compare the width of the 90% credible intervals for the chirp mass (left) and the mass ratio (right). The histograms show the ratio of the interval width after and before applying DeepClean, i.e., V1:DC / V1:ORG, implying that a value below 1 should indicate a reduction in uncertainty.
Even in cases where DeepClean effectively performs no subtraction, i.e., when the predicted noise is negligible, there is still an inherent randomness in the predicted noise realization. This is due to both the stochastic nature of the targeted noise coupling and the finite-duration nature of the regression analysis. This adds an additional layer of uncertainty to the cleaned data, and can reflect in the analysis data where the BBH signals are injected, as well as the PSD estimated from neighboring segments. These fluctuations can, in turn, cause small variations in the posterior widths. In an idealized scenario with no real subtraction, one might expect the error bar ratios (after vs. before DeepClean) to form a delta function at 1. However, in practice, due to the reasons outlined above, the distribution instead peaks around 1 but with spread on either side.
This is precisely what we observe for the mass ratio in the right panel of Figure 13: the histogram of error bar ratios is centered near 1 with roughly symmetric fluctuations. This behavior is consistent with a scenario where DeepClean does not yield significant improvement for this parameter and should not be interpreted as evidence of any adverse effect from the subtraction. In contrast, for the chirp mass—the most sensitive parameter in an inspiraling compact binary merger—we observe a clear shift toward ratios less than 1, indicating a reduction in uncertainty. In summary, both parameters exhibit a spread in their error bar ratio distributions due to the randomness in the noise predictions, with chirp mass showing an overall improvement and mass ratio reflecting neutral behavior.
Figure 13: Histogram of error bar ratios (DeepClean/original) for chirp mass (left) and mass ratio (right). Chirp mass shows a clear reduction in uncertainty, while mass ratio reflects neutral behavior with symmetric fluctuations around unity.
V Summary and Conclusions
In this paper, we applied the DeepClean framework to the Virgo detector data and demonstrated the successful subtraction of non-linear noise couplings across a wide frequency range. Between 15–415 Hz, we divided the strain data into 13 non-overlapping frequency bands and performed noise regression sequentially across them. Through this study, we demonstrate the effectiveness of non-linear noise regression in Virgo using DeepClean, resulting in improved detector sensitivity and enhanced detection prospects for GW signals. We have noted up to a 1.3 Mpc improvement in the BNS inspiral range ($\sim$2.5% gain).
Importantly, our validation tests show that the astrophysical signals are not adversely affected by the regression process, as confirmed through parameter estimation consistency checks. Additionally, we observed an increase in matched-filter SNR for multiple injected signals, which revealed an average increase of 1.7% in the recovered SNR for injected BBH. These results suggest that such regression methods can contribute meaningfully to future observing campaigns.
While DeepClean has been previously demonstrated with the two LIGO detectors, showcasing its capability for non-linear noise regression in Virgo—a detector with distinct design features—carries importance as a proof for their generalizability beyond their original deployment. Specifically, it validates the broader applicability of such tools and serves as a test for several Virgo-specific aspects that may challenge noise-regression efforts, such as channel availability, coupling stability, and noise behavior.
One of the longstanding challenges in noise regression has been the presence of a large number of witness channels, many of which exhibit non-negligible coherence with the strain channel. Traditional filtering techniques often struggle to process more than a few tens of such channels due to analytical and computational complexity. In contrast, deep-learning-based approaches such as DeepClean offer a scalable alternative. In this work, even within our band-split strategy, we have one case where 122 witness channels were used simultaneously in a single-pass training (see Fig. 7), and this configuration outperformed the multi-layer strategy that grouped channels by category within that particular band. This not only confirms DeepClean’s capacity to handle large numbers of channels within memory constraints, but also highlights the presence of meaningful cross-category couplings that are best captured when training is performed jointly. While there are practical memory limits, our analysis demonstrates a level of channel multiplicity far beyond what traditional methods can handle.
While our initial selection of witness channels was based on Brute Force Coherence—primarily identifying linear noise couplings—DeepClean inherently enables the removal of non-linear and non-stationary noise couplings that may exist between those channels, as we have shown in Section III.5. However, identifying a complete set of witness channels capable of capturing all such non-linear couplings goes beyond what Brute Force Coherence can identify. This would require a systematic exploration and validation of a broader set of witness channels, extending beyond linear coherence. Addressing this challenge is an open problem and is beyond the scope of the present work.
In the future, there are a number of challenges we plan to tackle:
  • Online analysis: All analysis in this paper was performed offline. However, online processing will be critical in O5 to leverage noise improvements in real time and enhance pre-merger GW detection. To support this, we plan to transition from the sequential band-wise cleaning strategy used in this work—where the output of one band feeds into the next—to a parallel band-wise architecture. This will allow training and cleaning across multiple frequency bands simultaneously, significantly improving computational efficiency and reducing latency. The core algorithmic components for this parallelization have already been developed and tested on LIGO data, and can be readily adapted for Virgo. We also iterate that this parallelization occurs across frequency bands (not across witness channels), and as long as these bands remain non-overlapping, there is no conflict even if some channels contribute to multiple bands.
  • High computational demands: Capturing the 150 Hz peak and its sidebands using ASC witness sensors required considerable computational resources, often requiring trade-offs between data length and resolution. We plan to explore lighter versions of the architecture for simpler forms of coupling, as well as alternative training strategies to reduce memory usage. For instance, all relevant noise couplings in this study are below 500 Hz, suggesting that a sampling rate of 1 kHz or lower may be sufficient. Nevertheless, our current pipeline operates at 4096 Hz to match astrophysical data needs. With careful tuning, lower sampling rates in the training phase may help optimize performance without compromising scientific goals.